OpenStudy (anonymous):

Hi. Write the complex number √2 - i in the trigonometric form r(cosθ + i sinθ), with θ in the interval [0°, 360°). Thanks in advance!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

root3(cos(315 + isin315)

OpenStudy (anonymous):

you need r and theta \[r=\sqrt{a^2+b^2}=\sqrt{\sqrt{2}^2+1}=\sqrt{3}\]\]

OpenStudy (anonymous):

to find theta you know what \[tan(\theta)=\frac{b}{a}=\frac{-1}{\sqrt{2}}\]

OpenStudy (anonymous):

you are in quadrant IV since you are to the right two and down one. so as himi1618 said \[\theta=315\]

OpenStudy (anonymous):

\[\sqrt{3}\]*(cos(315) + isin)315)) to find the quadrant , observe the sign in the given complex number. sqrt(2)-i, so its (+,-) and hence in 4th quadrant.

OpenStudy (anonymous):

and you are done \[\sqrt{2}-i=\sqrt{3}(\cos(315)+isin(315))\]

OpenStudy (anonymous):

i have to add that it is somewhat objectionable to use degrees in these problems but if that is what it said, that is what you do

OpenStudy (anonymous):

OK thanks for your help. Just one question: should I just now use a calculator to find theta in degrees? These always confuse me...